OpenStudy (anonymous):

Which is approximately equal to the area of the circle? Use 3.14 for Pi

OpenStudy (anonymous):

\[\pi r^2\]

Parth (parthkohli):

Area is pi * r*r \(\Large \color{Black}{\Rightarrow 3.14r^2 }\)

OpenStudy (shubhamsrg):

none..area = pi (r^2) when r=1,,then the above ques comes into play for approximation,,3.14 is used.. but if you want to be more and more precise..use (pi) ..the more digits you use of pi,,the more exact result you get..<--but last step is useless!!

OpenStudy (anonymous):



Parth (parthkohli):

Then that'd be: \(\Large \color{Black}{\Rightarrow 3.14 \times 4 \times 4 }\)

OpenStudy (anonymous):

thats ur radius

OpenStudy (anonymous):

Parth are u sure that will give me my answer

OpenStudy (anonymous):

I vouch for @ParthKohli

Parth (parthkohli):

It is \(\approx\) 48.42

Parth (parthkohli):

Thank you @Kiki547 :) you're now a good lad on OpenStudy :D

OpenStudy (anonymous):

i am a girl u know right don't u mean lassie lol

OpenStudy (lgbasallote):

he's sexist...he calls everyone a guy

Parth (parthkohli):

lol i'm a bad boy xD

OpenStudy (unklerhaukus):

\[\pi\approx3.14,\qquad\qquad r=4\]\[A_{circle}=\pi r^2\approx3.14\times4^2=3.14\times16=48+1.6+0.64=50.24\]

Parth (parthkohli):

Where did that 1.6 come from?

OpenStudy (unklerhaukus):

\[0.1\times16=1.6\]

OpenStudy (unklerhaukus):

\[3.14\times16=3\times16+0.1\times16+0.04\times16\]